The PCB concentration in the air inside a typical house is 400 ng/m3 , and the ventilation rate is such that it takes, on average, 10 h to change the indoor air. If the indoor PCBs are coming from a leaking capacitor, how much leaks per year?

ANSWER IS 160MG PER YEAR, but need help reaching that solution
Answer #1

In order to solve this question, the volume of the house is also needed.

If we take the area for this typical house to be 185.80608 m2 and its average height to be 2.5 m. Then,

Total Volume of the House = 185.80608 m2 x 2.5 m = 464.5152 m3

Total amount of PCBs = 464.5152 m3 x 400 x 10-9 g/ m3 = 1.86 x 10-4

Thus the Flow = 1.86 x 10-4/10 hrs  = 1.858 x 10-5 g/hr = 0.1627 g/yr = 162.7 mg/yr